# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

### v2

@@ 6 Additional Analysis
InternVL2-40B Image 0.0 4.91 0.0 - 8.3 48.18 61.61 0.96 291.1
LLaVA-v1.6-34B Image 0.0 0.54 0.0 - 15.0 34.79 64.97 1.47 289.8
LLaVA-OneVision-7B Image 0.0 1.25 0.0 - 8.3 41.02 62.89 0.80 294.9
Table 4: Comparison of scores between different run types: using only VLM (which receives an image as input) and using a mixture of VLM with LLM (which receives a caption as input). Metrics improved with caption inputs are highlighted in bold. Base LLM refers to the language backbone underlying the corresponding VLM. Base LLM (R1) for InternVL2.5-38B denotes Deepseek-R1-Distill-Qwen-32B. Experiments without step counts all ended in failure due to reaching the maximum turn limit.
@@ 6.1 Repetitive Ineffective Action Patterns
To analyze why higher success rates and goal completion correlate with reduced average steps in Table 3 despite the expectation that more steps would be required for higher SR and GC, we observe that lower-performing models frequently repeat the same action in identical situations throughout their trajectories. Figure 7 shows that these redundant attempts are more prevalent in lower-performing models, including both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Consequently, they accumulate longer trajectories, which in turn leads to lower SPL scores.
Figure 7: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions) on $exp_{\text{base}}$ setting. The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
@@ 6.3 Replacing VLM with LLM
Socratic models methodology, as represented by [36], is an approach that uses image captioners to convert visual inputs into descriptive text inputs and then employs LLMs to perform multimodal inference. Since it is widely known that VLMs excel at visual perception but often struggle with complex visual reasoning [22, 14, 26], we aim to address this gap by adopting LLMs.
@@
Among open-source models, InternVL2.5-38B, which demonstrated performance comparable to GPT-4o-mini in the $exp_{\text{hint}}$ setting, showed a decrease of nearly half in success rate in both $exp_{\text{base}}$ and $exp_{\text{hint}}$ settings when LLM was adopted instead of VLM, with significant deterioration in all other metrics as well. In contrast, the remaining open-source models demonstrated an average x3.1 increase in SR. A particularly significant observation pertains to the case of InternVL2.5-38B; upon substituting the LLM Qwen-32B-Instruct with Deepseek-R1-Distill-Qwen-32B, which shares an identical model architecture but has undergone reasoning capability distillation from a state-of-the-art reasoning model (Deepseek-R1 [15]) and thus possesses remarkably enhanced reasoning capabilities, all evaluation metrics improved notably. These results suggest that: 1. Many VLMs struggle with complex reasoning in multimodal contexts—a critical capability for VisEscape, and 2. Integrating strong reasoning capabilities of LLMs with visual perception of VLMs can lead to enhanced multimodal reasoning capability.
@@ 7 Conclusion
8 Limitation
While VisEscape presents a valuable benchmark for testing the exploration-driven problem-solving abilities of multimodal agents, the specific task of room escapes may not fully generalize to real-world environments where agents encounter more complex situations and more unfamiliar objects. Additionally, since the action space of VisEscape is discrete rather than continuous, there is a need to convert the conclusions drawn from multiple modules into feasible continuous action sequences for real-world execution.
9 Acknowledgements
This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name: Development of multimodal UX evaluation platform technology for XR spatial responsive content optimization, Project Number: RS-2024-00361757).
Supplementary Material
@@ 9 Acknowledgements
  • In Appendix A, we provide results and analysis for additional experiments that were not covered in the main paper.
  • In Appendix B, we provide comprehensive details about our experimental setup and the prompt template used.
@@
  • In Appendix D, we present details about the conducted human study.
Appendix A. Additional Experiments and Analysis
A.1. Tendency of Brute-Force Approaches
@@
In Figure 7, we show that low-performing agents exhibited meaningless, repetitive action patterns. However, upon analyzing the trajectories of various agents, we found that, when attempting to solve visual quizzes, smaller models frequently resorted to brute-force strategies—randomly trying different passwords or sequences without grounding their actions in any meaningful clues. For example, for problems requiring agents to arrange a given set of numbers in the correct order, some agents repeatedly attempted various possible combinations. To verify whether this behavior constituted brute-force attempts, we directly annotated whether the reasoning process and the corresponding attempted-answer pairs indicated a brute-force approach or not. Specifically, we annotated actions according to the following rules: (1) If the reasoning was grounded in meaningful clues or observations, we did not annotate it as brute-force. (2) If the attempted answers involved repeatedly trying different arrangements or combinations without valid justification, then we annotated them as brute-force.
@@
Table 6 demonstrates that while Claude-3.5-Sonnet and GPT-4o typically grounded their puzzle-solving attempts in meaningful visual clues, GPT-4o-mini, InternVL2.5-38B and LLaVA-v1.6-34B frequently resorted to brute-force approaches. Additionally, models that made a greater effort to identify reasoning justifications for their answers exhibited higher success rates in actual experiments. This behavior partially explains why these lower-performing models consistently achieved lower SPL and step efficiency scores—they made more attempts and fail more frequently on quizzes.
@@
Table 6: Results for puzzle-solving. Brute-force attempt ratio denotes the ratio of brute-force attempts among all answers submitted for numeric or alphabetic lock puzzles. Solved quiz ratio denotes the ratio of correctly solved locks to the total number of locks in VisEscape under the $exp_{\text{base}}$ setting.
A.2. Effect of the Exploration Memory
According to Table 5, removing the Memory module notably reduced VisEscaper performance. Beyond performance metrics, we hypothesize that among the three components of Memory module, particularly the Exploration memory component, which tracks already explored versus unexplored scenes or objects, encourages agents to seek out unseen scenes or objects. To verify this, we conducted an additional ablation experiment in which we specifically removed the Exploration memory, while keeping the other two components (Structured spatial memory and Salient action memory) intact within the Memory module. For both settings, we measured the proportion of scenes observed by the agent from the set of essential observations (as determined by the oracle trajectory) required to successfully complete the game. Figure 9 shows that InternVL2.5-38B exhibited only a slight difference in exploration of unique scenes, whereas LLaVA-v1.6-34B shows a substantial decrease.
A.3. Visualization of Interactions in VisEscape
@@
A.4. Statistics for Image Captions
Since the Memory module uses image captions to construct and manage long-term memory, it is crucial to evaluate how accurately the VLM captures essential information about the current state of the scene. To evaluate captions, we define the ground-truth (GT) state as the complete set of all receptacles and items present in an image obtained from the internal game engine. We then parse the generated captions and examine whether they contain the information consistent with the GT state. Table 7 shows that captions generated by InternVL2.5-38B provide the most accurate information in the most concise way, even when compared to GPT-4o.
Models Accuracy(%) Avg. Length
@@ A.4. Statistics for Image Captions
A.5. Full Results of Table 4
@@
Appendix B. Experimental details
B.1. List of LLMs and VLMs Used
In the list below, we list the versions of the LLMs used in our experiments. For GPT and Claude models, we use the officially released versions, and for the others, we use the model versions available on Huggingface.
  • Claude-3.5-Sonnet  [3]:
    claude-3-5-sonnet-20241022
  • GPT-4o  [1]:
    gpt-4o-2024-08-06
  • GPT-4o-mini  [1]:
    gpt-4o-mini
  • LLaVA-v1.6-34B [23]:
    llava-hf/llava-v1.6-34b-hf
  • LLaVA-OneVision-Qwen2-7B [20]:
    llava-hf/llava-onevision-qwen2-7b-ov-hf
  • InternVL2-40B [6]:
    OpenGVLab/InternVL2-40B
  • InternVL2.5-38B [6]:
    OpenGVLab/InternVL2_5-38B
  • Qwen2-7B-Instruct [42]:
    Qwen/Qwen2-7B-Instruct
  • Qwen2.5-32B-Instruct [43]:
    Qwen/Qwen2.5-32B-Instruct
  • Yi-34B [46]:
    NousResearch/Nous-Hermes-2-Yi-34B
  • DeepSeek-R1-Distill-Qwen-32B [15]:
    deepseek-ai/DeepSeek-R1-Distill-
    Qwen-32B
B.2. VLMs and Corresponding Base LLMs from Table 4
As shown in Table 4, we used only the LLM that served as backbone language model of corresponding VLM for controlled experiments. Base LLMs corresponding to each VLMS used in Table 4 is shown in Table 9333The documents specifying the LLMs selected as language backbone models are as follows: InternVL2.5-38B: https://huggingface.co/OpenGVLab/InternVL2_5-38B InternVL2-40B: https://huggingface.co/OpenGVLab/InternVL2-40B LLaVA-v1.6-34B: https://huggingface.co/llava-hf/llava-v1.6-34b-hf LLaVA-OneVision-7B: https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf
Model Type Model Name
VLM InternVL2.5-38B
Base LLM Qwen2.5-32B-Instruct
Base LLM (R1) DeepSeek-R1-Distill-Qwen-32B
VLM InternVL2-40B
Base LLM Yi-34B
VLM LLaVA-v1.6-34B
Base LLM Yi-34B
VLM LLaVA-OneVision-7B
Base LLM Qwen2-7B-Instruct
Table 9: VLMs and their corresponding base LLMs used in experiments on Table 4.
B.3. Computational Resources and Environments
All results reported in this paper are aggregated from three different runs conducted across 20 rooms. The time required to complete each room ranged from as little as 10 minutes (for models under 10B parameters, which often terminated early due to maximum turn limit) to as long as 90 minutes (for Deepseek-R1-Distill-Qwen-32B, which required considerable time for reasoning). For closed-source models, GPT-4o and Claude-3.5-Sonnet cost approximately \$2–3 per room for three separate experiments, totaling about \$50 for a complete run.
B.4. Prompt templates
Table 1011121314 shows experiment prompts.

### v3

@@ 8 Additional Analysis
InternVL2-40B Image 0.0 4.91 - - 8.3 48.18 61.61 0.96 291.1
InternVL2-40B + Base LLM Caption 0.0 3.88 - - 32.0 68.00 48.22 4.39 250.6
LLaVA-v1.6-34B Image 0.0 0.54 - - 15.0 34.79 64.97 1.47 289.8
LLaVA-v1.6-34B + Base LLM Caption 0.0 4.14 - - 38.3 61.80 50.08 4.59 257.7
LLaVA-OneVision-7B Image 0.0 1.25 - - 8.3 41.02 62.89 0.80 294.9
LLaVA-OneVision-7B + Base LLM Caption 0.0 1.62 - - 25.0 58.93 70.70 2.40 287.3
Table 4: Comparative performance of results from MLLMs and Socratic method. Experiment is conducted by models with memory management and reasoning module. $exp_{\text{base}}$ and $exp_{\text{hint}}$ denote experiments conducted without hints and hint-guided experiments, respectively.
@@ 8.1 Repetitive Ineffective Action Patterns
To analyze why higher success rates and goal completion correlate with reduced average steps in Table LABEL:tab:main despite the expectation that more steps would be required for higher SR and GC, we observe that lower-performing models frequently repeat the same action in identical situations throughout their trajectories. Figure 10 shows that these redundant attempts are more prevalent in lower-performing models, including both $exp_{\text{base}}$ and $exp_{\text{hint}}$. Consequently, they accumulate longer trajectories, which in turn leads to lower SPL scores.
Figure 5: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions). The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
@@ 8.3 Replacing VLM with LLM
Socratic models methodology, as represented by shin2024socratic, is an approach that uses image captioners to convert visual inputs into descriptive text inputs and then employs LLMs to perform multimodal inference. Since it is widely known that VLMs excel at visual perception but often struggle with complex visual reasoning li2024enhancing; gan2022vision; liu2023matcr, we aim to address this gap by adopting LLMs.
@@
Among open-source models, InternVL2.5-38B, which demonstrated performance comparable to GPT-4o-mini in the $exp_{\text{hint}}$ setting, showed a decrease of nearly half in success rate in both $exp_{\text{base}}$ and $exp_{\text{hint}}$ settings when LLM was adopted instead of VLM, with significant deterioration in all other metrics as well. In contrast, the remaining open-source models demonstrated an average x3.1 increase in SR. A particularly significant observation pertains to the case of InternVL2.5-38B; upon substituting the LLM Qwen-32B-Instruct with Deepseek-R1-Distill-Qwen-32B, which shares an identical model architecture but has undergone reasoning capability distillation from a state-of-the-art reasoning model (Deepseek-R1 guo2025deepseek) and thus possesses remarkably enhanced reasoning capabilities, all evaluation metrics improved notably. These results suggest that: 1. Many VLMs struggle with complex reasoning in multimodal contexts—a critical capability for VisEscape, and 2. Integrating strong reasoning capabilities of LLMs with visual perception of VLMs can lead to enhanced multimodal reasoning capability.
@@ 9 Conclusion
10 Related Works
10.1 Interactive Environments for Evaluating LLM/VLM
Traditionally, text-based games have served as a primary benchmark for language models to explore given environments and learn behaviors in an interactive world cote2019textworld; hausknecht2020interactive; qian2024escapebench. More recently, text-based benchmarks have been used as a venue to evaluate various capabilities and tendencies of LLMs in dynamic situations, such as utilizing scientific knowledge wang2022scienceworld and ethical decision-making pan2023rewards.
Several studies have advanced LLM/VLM-based agents built on Minecraft wangvoyager; Odyssey2024; dong2024villageragent; guss2019minerl. In parallel, embodied agent simulators have facilitated evaluating language-driven agents interacting within visually grounded environments ai2thor; puig2018virtualhome; shridharalfworld; habitat19iccv.
Meanwhile, the GUI environment has recently been proposed as a benchmark deng2024multi; deng2023mind2web; xu2021grounding for evaluating VLMs’ interaction with visual content. This benchmark was designed to evaluate VLMs’ ability to perceive visual elements on web interfacesliuvisualwebbench; cheng2024seeclick, reason about their functionality shahbandeh2024naviqate, and execute actions to successfully complete given tasks shi2017world; furuta2023exposing; yao2022webshop; zhouwebarena.
10.2 LLM/VLM as an Action Decision Maker
Deploying LLM or VLM agents in real-world applications requires effective handling of dynamic and unpredictable environments mnih2013playing. Such environments are characterized by changing contexts, unexpected events, and continuous novel observations nagabandi2019learning; zhang2018adaptive. To successfully adapt, agents must actively self-evolve using mechanisms for adaptive reasoning and continuous learning zhou2024symbolic.
Recent studies have highlighted three critical capabilities for agent adaptation. First, accumulation of experience enables agents to remember previous interactions, particularly failures, and use these experiences to improve future decisions feng2024agile. Second, strategic exploration allows agents to intelligently choose actions based on past outcomes and external feedback, balancing the benefits of learning against potential risks zhu2023ghost. Third, knowledge abstraction helps agents form generalized high-level representations from raw observations, significantly improving their adaptability in diverse situations zhao2023expel.
To support these adaptation skills, we integrate three specialized modules into our agent architecture; Memory, Feedback, and ReAct yao2023react. Together, these modules strengthen agents’ reasoning, planning, and reflection capabilities, enabling effective performance in dynamic settings.
Appendix A Example Appendix
@@
  • In Appendix B, we provide details about our experimental setup and the prompt template used.
@@
  • In Appendix D, we provide results and analysis for additional experiments that were not covered in the main paper.
  • In Appendix  E, we provide additional details not extensively covered in the main paper.
  • In Appendix F, we briefly introduce and analyze works that address the topic of escape rooms.
  • In Appendix G, we present details about the human study conducted.
Appendix B Experimental details
B.1 Model List
In the list below, we list the versions of the LLMs used in our experiments. For GPT, Gemini, and Claude models, we use the officially released APIs, and for the others, we use the model versions available on Huggingface.
  • Claude-3.5-Sonnet  claude3:
    claude-3-5-sonnet-20241022
  • GPT-4o  achiam2023gpt:
    gpt-4o-2024-08-06
  • GPT-4o-mini  achiam2023gpt:
    gpt-4o-mini
  • gemini-1.5-Pro  team2024gemini:
    Gemini-1.5-Pro
  • gemini-2.0-Flash  google2024gemini:
    Gemini-2.0-Flash
  • Qwen2.5-VL-32B-Instruct bai2025qwen2_5vl:
    Qwen/Qwen2.5-VL-32B-Instruct
  • LLaVA-v1.6-34B liu2023visual:
    llava-hf/llava-v1.6-34b-hf
  • LLaVA-OneVision-Qwen2-7B li2024llava:
    llava-hf/llava-onevision-qwen2-7b-ov-hf
  • InternVL2-40B chen2024expanding:
    OpenGVLab/InternVL2-40B
  • InternVL2.5-38B chen2024expanding:
    OpenGVLab/InternVL2_5-38B
  • Qwen2-7B-Instruct yang2024qwen2:
    Qwen/Qwen2-7B-Instruct
  • Qwen2.5-32B-Instruct yang2024qwen2_5:
    Qwen/Qwen2.5-32B-Instruct
  • Yi-34B young2024yi:
    NousResearch/Nous-Hermes-2-Yi-34B
  • DeepSeek-R1-Distill-Qwen-32B guo2025deepseek:
    deepseek-ai/DeepSeek-R1-Distill-
    Qwen-32B
B.2 MLLMs and Corresponding Backbone LLMs
As shown in Table 4, we used only the LLM that served as backbone language model of corresponding MLLM for controlled experiments. Base LLMs corresponding to each MLLMs used in Table 4 are shown in Table 6222The documents specifying the LLMs selected as language backbone models are as follows: InternVL2.5-38B:
https://huggingface.co/OpenGVLab/InternVL2_5-38B
InternVL2-40B:
https://huggingface.co/OpenGVLab/InternVL2-40B
LLaVA-v1.6-34B:
https://huggingface.co/llava-hf/llava-v1.6-34b-hf
LLaVA-OneVision-7B:
https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf
Model Type Model Name
VLM InternVL2.5-38B
Base LLM Qwen2.5-32B-Instruct
Base LLM (R1) DeepSeek-R1-Distill-Qwen-32B
VLM InternVL2-40B
Base LLM Yi-34B
VLM LLaVA-v1.6-34B
Base LLM Yi-34B
VLM LLaVA-OneVision-7B
Base LLM Qwen2-7B-Instruct
Table 6: MLLMs and their corresponding base LLMs.
B.3 Computational Resources and Environments
All results reported in this paper are aggregated from three different runs conducted in 20 rooms. The time required to complete each room ranged from as little as 10 minutes (for models under 10B parameters, which often terminated early due to the maximum turn limit) to as long as 90 minutes (for Deepseek-R1-Distill-Qwen-32B, which required considerable time for reasoning). For closed-source models, GPT-4o and Claude-3.5-Sonnet cost approximately \$2–3 per room for three separate experiments, totaling about \$50 for a complete run.
B.4 Hints
To facilitate the evaluation of the agent’s progression along its path, all game logic in VisEscape is designed to be sequential. Therefore, the next checkpoint the agent must achieve after its current one is predetermined, as is the hint to guide it there. Hint messages are provided only during Hint-guided experiments if the agent gets stuck for 30 or more steps (i.e., shows no change in checkpoint progression).
Here are two examples of hint messages.
  • Object interaction: “Use key to unlock the wardrobe at the south wall.”
  • Answer for lock: “Solve the numberlock in the safe at the east wall. The answer is 8056.”
B.5 Prompt templates
Table 1415161718 shows experiment prompts.
Appendix C Dataset Construction Process
C.1 Asset Creation Using Trimble SketchUp
We use Trimble SketchUp to create object assets (receptacles and items) for the room escape dataset. Receptacles are designed with predefined states, such as open/closed, locked/unlocked, and pushed/pulled, enabling agent interactions. To accommodate scene dynamics, we model these predefined states and generate corresponding receptacles to reflect those states. The visualized examples of receptacles and items are shown in Figure 6.
Figure 7: Comparison of applying visual rendering using Chaos Enscape. The left shows the view in Autodesk Revit before using Chaos Enscape, and the right shows the view in Chaos Enscape.
C.2 Scene Creation Using Autodesk Revit
  • Receptacle View: A close-up view of a receptacle accessed through the “inspect” action from the Wall View. It changes to reflect the states of receptacles.
C.3 Enhancing Photo-realism Using Chaos Enscape
To enhance the realism of the dataset, we use Chaos Enscape, a rendering tool that processes scene data generated in Autodesk Revit. Chaos Enscape improves the visual fidelity of the scene, providing photorealistic object representations. This increased realism facilitates effective agent reasoning and interaction within the environment, enabling accurate simulation and analysis of escape room scenarios. Refer to Figure 7 to check the before and after application of Chaos Enscape. Visualization of other rooms is in Figure 22.
C.4 Game Logic Design
We used OpenAI’s web interface GPT models to design efficient and automated game logic. The authors then manually reviewed the generated game logic design to ensure error-free gameplay. The prompt used for game logic design is in Table 19.
Appendix D Additional Experiments and Analysis
D.1 Tendency of Brute-Force Approaches
@@
We found that when attempting to solve visual quizzes, smaller models frequently resorted to brute-force strategies—randomly trying different passwords or sequences without grounding their actions in any meaningful clues. For example, for problems requiring agents to arrange a given set of numbers in the correct order, some agents repeatedly attempted various possible combinations. To verify whether this behavior constituted brute-force attempts, we directly annotated whether the reasoning process and the corresponding attempted-answer pairs indicated a brute-force approach or not. Specifically, we annotated actions according to the following rules: (1) If the reasoning was grounded in meaningful clues or observations, we did not annotate it as brute-force. (2) If the attempted answers involved repeatedly trying different arrangements or combinations without valid justification, then we annotated them as brute-force.
@@
Table 7 demonstrates that while Claude-3.5-Sonnet and GPT-4o typically grounded their puzzle-solving attempts in meaningful visual clues, GPT-4o-mini, InternVL2.5-38B, and LLaVA-v1.6-34B frequently resorted to brute-force approaches. Additionally, models that made a greater effort to identify reasoning justifications for their answers exhibited higher success rates in actual experiments. This behavior partially explains why these lower-performing models consistently achieved lower SPL and step efficiency scores—they made more attempts and failed more frequently on quizzes.
@@
Table 7: Results for puzzle-solving. Brute-force attempt ratio denotes the ratio of brute-force attempts among all answers submitted for numeric or alphabetic lock puzzles. Solved quiz ratio denotes the ratio of correctly solved locks to the total number of locks in VisEscape.
D.2 Reasoning and Visual quizzes
In VisEscape, all visual quizzes require associative thinking abilities. Analyzing game trajectories of diverse models, we find that most failures were on visual quizzes. So, we investigated whether these were problems they inherently couldn’t solve, or if they failed because they couldn’t properly associate the key clues within the escape room environment. For this, we reformed them as multi-image Visual Question Answering (VQA) problems. By providing the model with both the clue image and the target lock image, we inform them that the two images are related to each other, and then have them solve the problem. Used prompts are in Table 20.
Furthermore, we compare pass@1 and pass@10 to determine if models could correctly solve the task via iterative reasoning, even when their initial attempt failed. To support this repeated hypothesizing and testing, the model’s previously generated reasoning sentence for an incorrect answer, along with the incorrect answer itself, was fed back as input for the subsequent attempt.
We selected one visual quiz from each room and then evaluated the performance of each model. Table 8 shows the number of correct answers by each model on these 20 visual quizzes. Results show that open-source models generally demonstrate lower performance (Pass@1 without reasoning) on these visual quizzes compared to frontier models. Also, Reasoning tends to improve the number of correctly solved quizzes across all models.
While frontier models like GPT-4o and Claude-3.5-Sonnet show substantial performance gains with reasoning, the extent of this improvement varies, with some open-source models also showing notable gains, particularly when more attempts are allowed (e.g., InternVL2.5-38B shows a +5 improvement in Pass@10 with reasoning).
Also, iterative reasoning, as indicated by the general increase in scores from Pass@1 to Pass@10 under reasoning conditions, leads to a higher number of correct answers. This suggests that an iterative process of hypothesis formulation and testing enables the models to reach the correct solutions.
Model Pass@1 Pass@10
No Reason Reason No Reason Reason
[…]
Table 8: Comparison of pass@1 and pass@10 scores for models: performance with and without reasoning.
D.3 Effect of the Exploration Memory
We hypothesize that among the three components of memory management module, particularly the Exploration memory component, which tracks already explored versus unexplored scenes or objects, encourages agents to seek out unseen scenes or objects. To verify this, we conducted an additional ablation experiment in which we specifically removed the Exploration memory, while keeping the other two components (Structured spatial memory and Salient action memory) intact within the Memory module. For both settings, we measured the proportion of scenes observed by the agent from the set of essential observations (as determined by the oracle trajectory) required to successfully complete the game. Figure 8 shows that InternVL2.5-38B exhibited only a slight difference in exploration of unique scenes, whereas LLaVA-v1.6-34B shows a substantial decrease.
@@ D.3 Effect of the Exploration Memory
D.4 Comparison on Progress
Figure 9 illustrates the progression of goal completion across all rooms during the first 100 steps. Notably, while Claude-3.5-Sonnet performs best overall, it exhibits a distinctly accelerated trajectory after approximately 30-40 steps. This acceleration indicates that strong models better manage accumulated room-state changes, effectively leveraging previous observations to make strategic and efficient decisions later in the trajectory.
Figure 9: Average GC (Goal Completion) progression across all rooms and experiments during the first 100 steps. The X-axis represents the number of steps, and the Y-axis represents the goal completion at each step. Solid lines represent results from hint-guided experiment, while dotted lines represent those without any hints.
Additionally, models exhibiting rapid progress (Claude-3.5-Sonnet and GPT-4o) used fewer hints compared to other models (low HCR in Table LABEL:tab:main_hint). This indicates that rapid progress by high-performing models does not stem from heavy reliance on hints; instead, they use hints selectively only to resolve specific obstacles and, once past these obstacles, accelerate progress by effectively leveraging information accumulated up to that point.
D.5 Repetition
To analyze the performance difference between frontier models and open-source models, we analyzed repetitive actions (repeating the same action in identical situations) exhibited by each model within their trajectories. We observe that lower-performing models frequently repeat actions they have already taken throughout their trajectories, getting stuck in repetitive cycles. Figure 10 shows that these redundant attempts are more prevalent in lower-performing models.
Figure 10: The ratio of actions attempted multiple times (two or more repetitions) within a single trajectory, distributed across the frequency of attempt counts (ranging from 2 to 10+ repetitions). The X-axis represents the number of repeated attempts, and the Y-axis shows the ratio of these attempts out of all actions.
D.6 Statistics for Image Captions
Since the Memory module uses image captions to construct and manage long-term memory, it is crucial to evaluate how accurately the MLLM captures essential information about the current state of the scene. To evaluate captions, we define the ground-truth (GT) state as the complete set of all receptacles and items present in an image obtained from the internal game engine. We then parse the generated captions and examine whether they contain the information consistent with the GT state. Table 9 shows that captions generated by InternVL2.5-38B provide the most accurate information in the most concise way, even when compared to GPT-4o.
Model Accuracy(%) Avg. Length
@@ D.6 Statistics for Image Captions
D.7 Full Results of LLM Experiment
@@
Model $exp_{\text{base}}$ $exp_{\text{hint}}$
Step(Mean.) Step(Std.) Step(Mean.) Step(Std.)
[…]
Table 11: Mean and standard deviation of gameplay steps for each model. The standard deviation was calculated using only successful escape trajectories. For models that failed to escape entirely, the standard deviation is marked with "-".
D.8 Reasoning Decompsition
The proportions of reasoning components and their composition for other models are shown in Table 12 and Figure 11.
Figure 11: Proportion for each reasoning component for other models.
Appendix E Detailed Explanation
E.1 Reasoning Component
We have classified the reasoning processes that models undertake while solving problems in VisEscape as follows: Observation, Guess, Hypothesis Formation, Planning, and Recall. The classification criteria are as follows:
  • Observation: Visually perceiving, describing, or analyzing the current scene.
  • Guess: Making a guess or assumption based on priors or knowledge, without exact evidence acquired in the environment.
  • Hypothesis Formation: Forming a hypothesis or idea to test in the environment, based on any evidence acquired in the environment.
  • Planning: Establishing or formulating action plans within the environment.
  • Recall: Retrieving information from history - remembering the state/location of objects, actions previously performed, and other relevant details.
We extract the pre-action reasoning and action performed by each model at every game step from the experimental logs, which were run with an applied memory management and a reasoning module. And they are annotated using GPT-4o achiam2023gpt. We set temperature as 0.0 for reproduction. To ensure the reliability of our machine-based evaluation method, we sampled 100 instances for each of the categories classified by GPT-4o, totaling 500 samples. These sentences were then manually annotated by humans following the same procedure. Figure 12 shows a heatmap between human and GPT-4o annotations, suggesting that while the model’s classifications largely align with human annotations, the model frequently misclassified the Guess type as Hypothesis Formation.
Figure 12: Heatmap of agreement between GPT-4o and human annotations for reasoning type classification. In 85.6% of the instances, the category annotations from GPT-4o and humans were identical.
Model Composition Ratio Length
[…]
Table 12: Reasoning composition for other models.
E.2 Memory Management Module
E.3 Visualization of Interactions in VisEscape
E.4 Example of Visual Quizzes
Figure 14 shows two examples of visual quizzes in VisEscape.
Appendix F Related works on Escape Rooms
EscapeBench qian2024escapebench proposes a benchmark to evaluate creative reasoning in text-only environments, with a focus on tool use as a proxy for measuring creativity. IDEA he2024ideaenhancingrulelearning presents an agent that mimics the loop of human rule-learning by integrating multiple forms of reasoning to guide its actions, focusing on solving text-based puzzles. MM-Escape wang2025multimodallargelanguagemodels aims to extend the escape room paradigm by building an embodied environment for the evaluation of MLLM agents. Departing from embodied setups, our visually grounded benchmark blends tool use with puzzle-solving to create a more flexible yet challenging reasoning framework for evaluating MLLMs.
Appendix G Human Study
Room SR(%) GC(%) Steps Duration(s)
1 75.0 92.5 41.8 231.3
2 100.0 100.0 67.3 319.5
3 50.0 87.5 71.1 319.5
4 75.0 93.8 112.0 447.0
5 100.0 100.0 77.3 274.5
6 100.0 100.0 37.0 117.0
7 100.0 100.0 46.8 225.0
8 100.0 100.0 53.0 567.3
9 75.0 90.0 52.0 567.3
10 75.0 91.7 54.0 204.5
11 50.0 79.2 37.8 231.3
12 75.0 92.9 31.3 144.0
13 100.0 100.0 41.5 199.8
14 100.0 100.0 57.3 468.3
15 100.0 100.0 36.0 151.0
16 100.0 100.0 40.8 106.0
17 25.0 96.9 63.3 488.3
18 100.0 100.0 40.5 399.3
19 75.0 91.7 43.8 378.5
20 75.0 100.0 50.8 530.3
Average 82.5 95.8 52.8 318.5
To compare the performance of AI models and human performance on escape room games, we conducted a human study. We recruited a total of 20 participants, including both individuals familiar and those unfamiliar with escape room games, and each was instructed to complete tasks in four different rooms. Figures 23 illustrate the Gradio interface used to provide the UI for playing VisEscape.
Table 13 presents the results of the human evaluation for 20 rooms, including the success rate, goal completion, number of steps, and time duration for experiments. Participants achieved high success rates in most rooms. We paid them based on the minimum hourly wage.